# OmniSage: Large Scale, Multi-Entity Heterogeneous Graph Representation Learning

### v1

@@ 3.2. Embedder Architecture
Given a node $u$ of type $t$ and its local neighborhood $N(u,\mathcal{H})$, the features of $u$ and $v\in N(u,\mathcal{H})$ are processed through a MLP and concatenated to form a sequence. This sequence is passed through transformer encoder layers with multi-head self-attention. The output from the last layer corresponding to a learned CLS token is processed through an MLP, followed by $L_{2}$ normalization to yield the final embedding, denoted as $f_{t}(u)$.
@@ 3.3.2. Entity-Feature Task
For each input feature type $x$, we initialize a multi-layer perceptron (MLP) encoder $g_{x}(u)$ as the learned encoding of the feature with type $x$. Similar to the entity-entity task, the entity-feature contrastive objective optimizes a sampled softmax loss across a dataset $\mathcal{D}_{\mathrm{feat}}$ containing nodes $\mathbf{u}$ of type $\mathbf{t_{u}}$ and a set of negative nodes $\mathbf{N}$ sampled from the same type for each $\mathbf{u}$. We reuse the softmax loss in Eq. 1, where the query is the feature encoding $g_{x}(\mathbf{u})$ for $x\in\mathcal{X}_{\mathbf{u}}$, the set of all feature types of node $u$, and the positive and negatives are the embedding of $\mathbf{u}$ and $\mathbf{N}$ respectively. The entity-feature task loss is
@@ 3.3.3. User-Entity Task
Following Pancha et al. (2022), we employ two sequence level objectives, namely the next action loss and the future action loss. The next action loss simply optimizes the user embedding to predict the next entity embedding, i.e., the query is a user embedding $\mathbf{U}_{t}$ at time $t$ and the matching positive is the next entity embedding $\mathbf{P}_{t+1}$. Given a sequence dataset $\mathcal{D}_{\mathrm{seq}}$, we sample in-batch negatives $\mathbf{N}$ for each batch of sequences. The next action loss is

### v2

@@ 3.2. Embedder Architecture
Given a node $u$ of type $t$ and its local neighborhood $N(u,\mathcal{H})$, the features of $u$ and $v\in N(u,\mathcal{H})$ (i.e., all neighbour features) are processed through a MLP and concatenated to form a sequence. This sequence is passed through transformer encoder layers with multi-head self-attention. The output from the last layer corresponding to a learned CLS token is processed through an MLP, followed by $L_{2}$ normalization to yield the final embedding, denoted as $f_{t}(u)$.
@@ 3.3.2. Entity-Feature Task
For each input feature type $x$, we initialize a multi-layer perceptron (MLP) encoder $g_{x}(u)$ as the learned encoding of the feature with type $x$. Similar to the entity-entity task, the entity-feature contrastive objective optimizes a sampled softmax loss across a dataset $\mathcal{D}_{\mathrm{feat}}$ containing nodes $\mathbf{u}$ of type $\mathbf{t_{u}}$ and a set of negative nodes $\mathbf{N}$ sampled from the same type for each $\mathbf{u}$. We reuse the softmax loss in Eq. 1, where the query is the feature encoding $g_{x}(\mathbf{u})$ for $x\in\mathcal{X}_{\mathbf{u}}$, the set of all feature types of node $u$, and the positive and negatives are the embedding of $\mathbf{u}$ and $\mathbf{N}$ respectively, as before. The entity-feature task loss is
@@ 3.3.3. User-Entity Task
Following Pancha et al. (2022), we employ two sequence level objectives, namely the next action loss and the future action loss. The next action loss simply optimizes the user embedding to predict the next entity embedding, i.e., the query is a user embedding $\mathbf{U}_{t}$ at time $t$ and the matching positive is the next entity embedding $\mathbf{P}_{t+1}$. Given a sequence dataset $\mathcal{D}_{\mathrm{seq}}$, we sample in-batch negatives $\mathbf{N}$ for each batch of sequences, again, as before. The next action loss is
@@ A.2.3. Visualization
A.2.4. Aggregation Technique Evaluation
Though we leverage a simple, out-of-the-box multi-head attention-based aggregation technique across our neighbourhood (and our anchor node’s features), we compare to some existing traditional aggregation techniques to ensure that our technique is optimal. For this purpose, we leverage Graph Attention Networks (GAT), which are reliable and perform strongly across a wide range of use cases (Veličković et al., 2017).
Due to computational restrictions, we limit our evaluation of GAT to 1-hop and 2-hop neighbours across the entity-entity task. We compare the performance of GAT with the proposed transformer-based aggregation used in OmniSage. We show the results in Table 11.
Table 11. Comparison of GAT in offline recall for aggregating neighbour features, relative to the transformer-based approach used in OmniSage.
Objectives Entity-Entity (lift, recall @ 10)
GAT (1-hop) -6.0%
GAT (2-hop) -8.5%
Though the GAT performs reasonably well with only the 1-hop neighbourhood, it performs even worse with the 2-hop neighbourhood, indicating that it is unable to adequately aggregate information from farther neighbours of the anchor node. On the other hand, OmniSage is able to aggregate nodes from across the entire neighbourhood, demonstrating superior performance compared to GAT across the board.
A.2.5. Ablating Negative Sampling Techniques
Following Pancha et al. (2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function applied for all our tasks. Though we effectively end up with similar conclusions as Pancha et al. (2022), we demonstrate these findings to justify our design choices.
Given that in-batch negatives are added with any additional processing cost (i.e., they are simply from the batch), we focus primarily on the effect of random negatives throughout this analysis. To ablate its effect, we examine the performance on the user-entity task with varying number of random negatives in Table 12. Clearly, it demonstrates that the presence of random negatives is crucial for performance, and additionally, scaling negatives has a positive effect (though not as positive, indicating a potential plateau).
Table 12. Lift in offline recall over using no random negatives in user-entity task.
# random negatives User-Entity (lift, recall @ 10)
1K +31.2%
10K +43.5%
Additionally, we discover that adding in-batch negatives without random negatives drops recall metrics (across a random index), suggesting that a random corpus of negatives during training is necessary for optimal retrieval on such a corpus at evaluation.
A.2.6. Additional Neighbour Sampling Analysis
To examine the performance scaling across the number of neighbours used, we halve the number of neighbours within each edge and node type selected for the top-k within the random walks. In doing so, we filter out the least relevant neighbours and provide less representational information. Does this affect the performance across any of our tasks? Note that given that the neighbourhood definition has changed, we do not evaluate the entity-feature objective.
We show the percentage lift across the entity-entity and user-entity tasks in Table 13. Though the margins are insignificant for the entity-entity task, there are reasonable lifts in the offline metrics for the user-entity tasks, showing value for additional neighbourhood representation.
Table 13. Lift in offline recall with halved set of top-k neighbours selected in random walks over default OmniSage configuration (i.e., as presented in the main text).
Task Lift in recall @ 10
Entity-Entity -0.1%
User-Entity -1.4%
Though some of the metrics are neutral with others moderately improved, we choose to keep the neighbourhood as large as possible for the purposes of our entity-feature task, motivated by the goal to represent as much of the heterogeneous graph neighbourhood as possible to lend the embedding the most representational power.
A.3. Implementation Details
To enable reproducibility, we describe several implementation details of OmniSage, including more details about the embedders used to produce the representations, the architecture for each of the tasks, and the specific hyperparameters and configuration. In addition to this, we are planning to open-source relevant segments of the model code to enable reproducibility and exploration from the wider research and industry community.
A.3.1. Embedders
Provided the features for the anchor node and its neighbours’ features, we pass these into a small 1-layer transformer aggregator with 12 heads and an embedding dimension of 768, with an MLP dimension of 3072. In order to aggregate the representations across the “tokens” (in this case, the concatenation of anchor node features and neighbour features), we leverage a CLS-style token at the beginning of the sequence. Consequently, we apply no (causal) masking to the sequence. The output of the CLS token is passed into a small MLP (denoted as the transformer head), which is subsequently normalized into the OmniSage embedding. Note that the described architecture is used both for the pin embedder and the board embedder, except the board embedder has a different sequence length owing to a different set of features.
A.3.2. Task Architecture
In this section, we describe the architecture of the tasks, which provide the necessary representational power to OmniSage.
Entity-Entity
There is no architecture specifically for the entity-entity task since it leverages the dot product of similar or dissimilar entities only. These similarities are passed directly to the sampled softmax loss function.
Entity-Feature
Though our decoder for the OmniSage embedding is the identity function, we encode the raw features using a simple, 4-layer MLP with a ReLU activation function. The MLP contains a hidden size of 1,024, and we do not perform normalization or layer norm.
User-Entity
To enable modeling of user sequences, we leverage a small causal transformer, with 4 transformer layers and 4 heads. The architecture is similar to Pancha et al. (2022), except we leverage the OmniSage embedding during training (with gradients through the embedders producing the embedding).
A.3.3. Hyperparameters
We provide a list of hyperparameters below for the configuration of various elements within OmniSage:
  • Embedder activation function: ReLU
  • Embedder transformer layers: 1
  • Embedder transformer heads: 12
  • Embedder transformer dimension: 768
  • Embedder MLP dimension: 3072
  • Feature projection activation function: ReLU
  • Feature projection layers: 4
  • Feature projection hidden dimension: 1024
  • User-entity transformer, activation function: GELU
  • User-entity transformer layers: 4
  • User-entity transformer heads: 4
  • User-entity transformer dimension: 512
  • Learning rate: 0.005
  • Learning rate scheduler: linear warmup, cosine annealing
  • Optimizer: AdamW
  • Precision: fp16